OpenStudy (anonymous):

9x+3y=6. What does the graph look like for this term ?

OpenStudy (anonymous):

Whoa, you, like, asked this question earlier this evening!

OpenStudy (anonymous):

No i didnt...this is totally different...

OpenStudy (anonymous):

\[9x+3y=6\\3x+y=2\\y=-3x+2\]

OpenStudy (anonymous):

bad references is correct, but an easier way to graph is to use the cover up method, 9x+3(0)=6 <--- You set either the x or y to zero and solve for the other. Answer is, x= 2/3 <--Where it crosses x axis 9(0)+3y=6 Answer is, y=2 <--Where it crosses y axis